Prior to June 30, a company has never had any treasury stock transactions. A company repurchased 100 shares of its common stock on June 30 for $40 per share. On July 20, it reissued 50 of these shares at $46 per share. On August 1, it reissued 20 of the shares at $38 per share. What is the balance in the Treasury Stock account on August 2
Answer #1

Treasury stock  as on June 30-----100*40=4000

Stock as July 20--------50*46=2300

Stock as on august 1--------20*38=760

Total treasury stock as on august 2 is -------------100+50+20=170

Total value of treasury stock as on august 2 is (100*40)+(50*46)+(20*38)=4000+2300+760

=7060